# Contextual Representation Anchor Network to Alleviate Selection Bias in Few-Shot Drug Discovery

### v1

@@ Abstract
In the drug discovery process, the low success rate of drug candidate screening often leads to insufficient labeled data, causing the few-shot learning problem in molecular property prediction. Existing methods for few-shot molecular property prediction overlook the sample selection bias, which arises from non-random sample selection in chemical experiments. This bias in data representativeness leads to suboptimal performance. To overcome this challenge, we present a novel method named contextual representation anchor Network (CRA), where an anchor refers to a cluster center of the representations of molecules and serves as a bridge to transfer enriched contextual knowledge into molecular representations and enhance their expressiveness. CRA introduces a dual-augmentation mechanism that includes context augmentation, which dynamically retrieves analogous unlabeled molecules and captures their task-specific contextual knowledge to enhance the anchors, and anchor augmentation, which leverages the anchors to augment the molecular representations. We evaluate our approach on the MoleculeNet and FS-Mol benchmarks, as well as in domain transfer experiments. The results demonstrate that CRA outperforms the state-of-the-art by 2.60% and 3.28% in AUC and $\Delta$AUC-PR metrics, respectively, and exhibits superior generalization capabilities.
@@ I Introduction
  • We introduce a context augmentation module that dynamically retrieves analogous molecules and captures task-specific contextual knowledge to augment the contextual representation anchors. By continuously integrating feedbacks from similar molecules, this module refines the positioning and structural relationships of the anchors within the feature space, ensuring that they accurately capture the nuanced co-occurrence patterns and dependencies between molecular properties (Section IV-B1).
@@ III-B Multi-Head Attention Mechanism
$=\sum_{i=1}^{H}\operatorname{softmax}\left(\dfrac{\mathbf{Q}\mathbf{W}_{i}^{Q}(\mathbf{K}\mathbf{W}_{i}^{K})^{\top}}{\sqrt{d_{k}}}\right)\mathbf{V}\mathbf{W}_{i}^{V}\cdot\mathbf{W}_{i}^{O}$
@@
IV Method
In this section, we present a detailed overview of our Contextual Representation Anchor Network (CRA), with its overall architecture illustrated in Figure 3. First, we define the few-shot learning problem in the context of molecular property prediction and introduce the role of the attention mechanism. Then, we describe the architecture of our CRA model, followed by an explanation of the CRA algorithm. Finally, we detail the training and inference phases of our model.
@@ VI Conclusion
In this work, we proposed the Contextual Representation Anchor Network (CRA) to address the challenge of sample selection bias in few-shot molecular property prediction. By introducing a dual augmentation mechanism—comprising context augmentation and anchor augmentation—CRA dynamically incorporates task-specific contextual knowledge, enhancing the expressiveness of molecular representations. Our extensive evaluation on MoleculeNet and FS-Mol benchmarks, as well as domain transfer experiments on the Tox21 dataset, demonstrates that CRA consistently outperforms state-of-the-art methods in terms of both AUC and $\Delta$AUC-PR metrics. Furthermore, CRA’s ability to generalize across domains highlights its potential for robust application in drug discovery scenarios where labeled data is scarce. In future work, we aim to further explore the application of CRA in broader molecular and biochemical datasets, as well as to investigate the integration of additional contextual information to further improve its performance.

### v2

# Contextual Representation Anchor Network for Mitigating Selection Bias in Few-Shot Drug Discovery

@@ Abstract
In the drug discovery process, the low success rate of drug candidate screening often leads to insufficient labeled data, causing the few-shot learning problem in molecular property prediction. Existing methods for few-shot molecular property prediction overlook the sample selection bias, which arises from non-random sample selection in chemical experiments. This bias in data representativeness leads to suboptimal performance. To overcome this challenge, we present a novel method named contextual representation anchor network (CRA), where an anchor refers to a cluster center of the representations of molecules and serves as a bridge to transfer enriched contextual knowledge into molecular representations and enhance their expressiveness. CRA introduces a dual-augmentation mechanism that includes context augmentation, which dynamically retrieves analogous unlabeled molecules and captures their task-specific contextual knowledge to enhance the anchors, and anchor augmentation, which leverages the anchors to augment the molecular representations.
We evaluate our approach using the MoleculeNet and FS-Mol benchmarks, as well as through domain transfer experiments. The outcomes indicate that CRA surpasses current state-of-the-art methods by 2.60% in AUC and 3.28% in $\Delta$AUC-PR metrics, showcasing its exceptional generalization abilities.
@@ I Introduction
  • We present a context augmentation module that dynamically retrieves analogous molecules and captures task-specific contextual knowledge to augment the contextual representation anchors. By continuously integrating feedbacks from similar molecules, this module refines the positioning and structural relationships of the anchors within the feature space, ensuring that they accurately capture the nuanced co-occurrence patterns and dependencies between molecular properties (Section IV-B1).
@@ III-B Multi-Head Attention Mechanism
$=\sum_{i=1}^{H}\operatorname{softmax}\left(\dfrac{\mathbf{Q}\mathbf{W}_{i}^{Q}(\mathbf{K}\mathbf{W}_{i}^{K})^{\top}}{\sqrt{d_{k}}}\right)\mathbf{V}\mathbf{W}_{i}^{V}\cdot\mathbf{W}_{i}^{O},$
@@
IV Methodology
In this section, we present a detailed overview of our contextual representation anchor network (CRA), with its overall architecture illustrated in Figure 3. First, we define the few-shot learning problem in the context of molecular property prediction and introduce the role of the attention mechanism. Then, we describe the architecture of our CRA model, followed by an explanation of the CRA algorithm. Finally, we provide a detailed introduction of the training and inference phases of our model.
@@ VI Conclusion
In this work, we present the contextual representation anchor network (CRA) to address the challenge of sample selection bias in few-shot molecular property prediction. By implementing a dual augmentation mechanism—consisting of context augmentation and anchor augmentation—CRA dynamically incorporates task-specific contextual knowledge, enhancing the expressiveness of molecular representations. Our extensive evaluations on the MoleculeNet and FS-Mol benchmarks, along with domain transfer experiments on the Tox21 dataset, demonstrate that CRA consistently outperforms state-of-the-art methods in both AUC and $\Delta$AUC-PR metrics. Furthermore, CRA’s ability to generalize across domains underscores its potential for robust applications in drug discovery, particularly where labeled data is limited. In future work, we aim to expand the application of CRA to broader molecular and biochemical datasets and to explore the integration of additional contextual information to enhance its performance further.